1. A 5.2kg piece of iron at 28.6°C absorbs 35.0kJ of energy.  What is its final temperature?
1
Expert's answer
2021-01-21T03:51:03-0500

Q=mc∆T. ∆T=Q/mc = 35000/5200 x 0.444 = 15,1593902.

28.6 + 15,1593902 = 43,7593902 °C